OpenStudy (anonymous):

the two sides of the triangle are 34 cm and 43 cm , if the included angle is 50 degree. Find the perimeter of the triangle?

OpenStudy (anonymous):

Since it's Side Angle Side, use the law of cosine and add all the sides. \[b^2=34^2+43^2-2(34)(43) \cos (50°)\]